General Meeting Summary 10/29/20

From Cuawiki
Jump to: navigation, search

PARTICIPANTS: Aaron Brown, Alexandre Camsonne, Bogdan Wojtsekhowski, Brad Sawatzky, Carlos Munoz-Camacho, Carlos Yero, Rolf Ent, Gabriel Niculescu, Hamlet Mkrtchyan, Julie Roche, Marie Boer, Steven Lassiter, Vardan Tadevosyan, Vladimir Berdnikov, Tanja Horn, Charles Hyde

SUMMARY NOTES

FRAME (Carlos)

  • Discussion of assembly status and temperature control system
  • proposed numbering scheme: start with bottom left and go up (view from back of detector)
  • Good progress with the assembly
  • getting ready to ship towards end of the year
  • some issues, e.g. pipe material and flexibility, with cooling identified and resolved


DSG UPDATES (Aaron)

  • All cables ordered
  • CAEN HV testing
  • all channels so far respond as expected
  • EPICS screen development ongoing
  • Control screens
  • numbering should be consistent with the one proposed by Carlos
  • Interlock system
  • temperature scanning system


MECHANICAL/DESIGN (Steven)

  • Making good progress - some of the ordered parts are starting to arrive
  • Paulo has been working with Walter about storing location


FAST ELECTRONICS GROUP UPDATE (Brad S.)

  • Full update at next meeting by Chris Cuevas
  • Cables:
  • RG58 cables - vendor started fabrication (2200 cables), should be ready end of year
  • All HV equipment should be there
  • Work in SHMS hut
  • need two racks, one additional rack has to be moved in
  • Reflector shaping tool - waiting for update from Vladimir/Josh
  • Chiller unit placement - proposal: on beam left of SHMS, length of chiller lines is 8.2m (detector moving back and forth - need flexible lines), needs less shielding this way
  • Space for assembly: need sketch for space needed - Brad can make a first draft


NPS SOFTWARE (all)

  • shower clustering, recon done by Ho San - standalone
  • worked in Hall A framework
  • Action: upload shower clustering/recon codes to github, bring in Steve Wood to compile into Hall C ANa, test with some older calorimeter data
  • fADC250 analysis still missing, raw waveform (not currently available)
  • possible suitable codes have been used for SBS, decoding for fADC250 - could perhaps be recycled - there should be a github with that
  • Action: review and make available the fADC decoding
  • Performance of software
  • NPS has blocks than Hall A - consider developing performance checks


COLLABORATION MEETING (All)

  • Focus on discussions (identify missing pieces) rather than formal presentations
  • Proposal: two half days (February 1-2, 2021)


HIGH VOLTAGE DIVIDER

  • Action: check with Fernando on anode current test plans


NEXT MEETING: THURSDAY 12 NOVEMBER 2020 at 9:00AM ET